OpenStudy (anonymous):

Find x - 3y if x = 3 and y = -2.

OpenStudy (aroub):

3-3(-2) 9

OpenStudy (anonymous):

is it positive 9 or negative

OpenStudy (aroub):

positive

OpenStudy (anonymous):

(3)- 3(-2)= 3+6= +9

OpenStudy (anonymous):

thank u